OpenStudy (anonymous):

Will give medals and fan!! How many solutions are there to the following system of equations? 2x – y = 2 –x + 5y = 3 A. 1 B. 0 C. 2 D. infinitely many

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

Do you understand this @tblue1998

OpenStudy (anonymous):

HELP PLEASE!!

OpenStudy (anonymous):

Hi give me a second

OpenStudy (anonymous):

2x – y = 2 –x + 5y = 3 This is a system of equations. I will use substitution.

OpenStudy (anonymous):

ok thanks!!!!

OpenStudy (anonymous):

-x + 5y = 3 -x = 3 - 5y x = 5y - 3 2x -y = 2 (substitute the top part into here) 2(5y-3) - y = 2 10y - 6 - y = 2 9y = 8 y = 8/9

OpenStudy (anonymous):

x = 5y - 3 = 5(8/9) - 3 = 13/9 = 1.444444 repeating Since there is a solution of (13/9, 8/9), there is 1 solution

OpenStudy (anonymous):

Thanks can you help me with another one?

OpenStudy (anonymous):

The answer is A. Keep in mind, answer choice C is impossible; you can never have a system of equations with 2 solutions :) Sure!

OpenStudy (anonymous):

Solve the following system of equations. x + y = 3 2x + 2y = 6 A. no solution B. infinite number of solutions C. (0, 3) D. (1.5, 1.5) @Mathexpert

OpenStudy (anonymous):

x + y = 3 2x + 2y = 6 I will use combination this time. x + y = 3 2(x + y) = 2(3) 2x + 2y = 3 2x + 2y = 3 2x + 2y = 6 These are my two equations (the first one I just rewrote to fit my way of solving). Now subtract the equations. 0 = -3 This is not true; thus, there are no solutions. The answer is A

OpenStudy (anonymous):

thanks i have 3 more think you can help? @Mathexpert

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Solve the following system of equations by linear combination: 3x + 2y = –7 5x – 4y = –19 A. The solution is (–3, 1). B. The solution is (–4, 2). C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

@Mathexpert

OpenStudy (anonymous):

You are an absolute lifesavor! @Mathexpert

OpenStudy (anonymous):

3x + 2y = –7 5x – 4y = –19 Using combination again. Rewriting top equation. 3x + 2y = –7 2(3x + 2y) = 2(-7) 6x+4y = -14 6x+4y = -14 5x-4y = -19 Add both equations. 11x = -33 x = -3 Plug back x into one of the equations to get y. 3x + 2y = -7 3(-3) + 2y = -7 -9 + 2y = -7 2y = 2 y = 1 the solution is (-3,1), choice A

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Which is the graph of the system of equations? 2x + 3y = 12 4x + y = 8 A. ALGTests_L2_S1_U8_Q11_B.gif B. ALGTests_L2_S1_U8_Q11_D.gif C. ALGTests_L2_S1_U8_Q11_A.gif D. ALGTests_L2_S1_U8_Q11_C.gif

OpenStudy (anonymous):

I can't see the graphs unfortunately.

OpenStudy (anonymous):

Use mathway.com to graph them.

OpenStudy (anonymous):

ALGTests_L2_S1_U8_Q11_B.gif B. ALGTests_L2_S1_U8_Q11_D.gif C. Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y + 2 A. (y + 2) + y = 6 B. x = (x – 6) + 2 C. x + (y + 2) = 6 D. x + y = y + 2 D. ALGTests_L2_S1_U8_Q11_C.gif

OpenStudy (anonymous):

I posted the wrong question!!!!

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y + 2 A. (y + 2) + y = 6 B. x = (x – 6) + 2 C. x + (y + 2) = 6 D. x + y = y + 2

OpenStudy (anonymous):

x + y = 6 x = y + 2 Subsitute x = y + 2 into x + y = 6 (y + 2) + y = 6 The answer is A.

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (anonymous):